Let X be a continuous random variable. Prove that: P(21-; < X < xạ) = 1 - a.
Answer #1

x will be between 1 and 0

between 1-a/2 and a/2 the range is : r = (1-a/2) - (a/2) = 1 - a

the probability of P(x in r) = r / total

total probability is always = 1

P(x in r) = r / total = (1-a) / 1 = 1-a